There have four questions!

6. Omega plc purchased a piece of equipment at a cost of £125,000. It is expected to have a useful economic life of 5 years and a scrap value of £25,000 at the end of its life. Omega uses straight line depreciation and a full year of depreciation is charged in the year of purchase. What would be the annual depreciation charge and the net book value (NBV) of the equipment at the end of the second year of its life? (a) (b) (c) (d) (e) Depreciation charge £25,000; NBV £100,000 Depreciation charge £25,000; NBV £75,000 Depreciation charge £20,000; NBV £85,000 Depreciation charge £20,000; NBV £80,000 Depreciation charge £30,000; NBV £65,000 7. Using the information in Question 6, what would be the depreciation charge in year 2, based on a reducing balance percentage of 30%? (a) £49,000 (b) £73,500 (c) £26,250 (d) £11,250 (e) £37,500 8. For a defined benefit pension plan, the investment risk lies with: (a) The pension fund manager (b) The pension fund members (the employees) (c) The pension fund sponsor (the firm) (d) The governments pension lifeboat scheme (e) s shared equally by member, manager and the sponsor 9. The Enterprise value (EV) of a firm is given by: (a) (b) (c) (d) (e) Book value of owners equity Market value of owners equity Book value of firms debt plus book value of owners equity Book value of firms debt plus market value of owners equity Market value of firms debt plus market value of owners equity
Answer #1

Ans to ques no. 6

Depreciation per year = (125,000-25,000) / 5 = 20,000 per year

Net book value at the end of year 2 = 125000 - (20000*2) = 85,000

Annual depreciation charged for year 2 = 20,000

So the correct option is (c) Depreciation charge 20000; NBV 85,000

Ans to ques no. 7

Depreciation charge to year 2 on reducing balance percentage of 30 %

Year 1 = 125000 - 30 % = 87,500 NBV

Year 2 Depreciation = 87,500 * 30% = 26,250

So corrct option is (c) i.e. 26,250

Ans to ques no. 8

Correct option is (c) i.e The pension fund sponsor ( the firm)

Ans to ques no. 9

Correct option is (e) i.e. Market value of firm's debt plus market value of owner's equity